A rectangle is 9ft longer than its wide, and its area is 360ft2. use quadratic equation to find its length and width
Answer by mananth(16949)

width =x
length = x+9
...
Area = 360 ft^2
x(x+9)=360
x^2+9x=360
x^2+9x-360=0
x^2+24x-15x-360=0
x(x+24)-15(x+24)=0
(x-15)(x+24)=0
So x = 15 . Ignore -24
width = 15 feet
length = 24 feet
...
m.ananth@hotmail.ca